# Towards Browser Controls to Protect Cookies from Malicious Extensions

### v2

@@ 5.1 BrowserOnly Attribute
Similarly, Figure 3(B) displays how the Network Service parses the server’s response. When the response is received (step B.1 in Figure 3(B)), the URLRequestHTTPJob forwards the response headers to extensions (B.2). Once the extensions finish, the headers are returned to the HTTPJob (B.3). At this point, any new cookies are added to the Cookie Monster (B.4) and the response is sent to the renderer for display (B.5). To hide BrowserOnly cookies in responses from extensions, we strip all Set-Cookie headers with the BrowserOnly attribute from the response on arrival (B.1). Then when the headers are returned from the extensions (B.3) we reattach the BrowserOnly headers to the response.
@@
For simplicity, this explanation omits some details of how the Network Service handles transactions and interacts with extensions. Requests pass through several classes between the renderer and the network. We focus on the URLRequestHTTPJob as all BrowserOnly (and Monitored) functionality was added to this class. Similarly, when building a network request, its headers are sent to extensions up to three times [59]. Protecting BrowserOnly cookies requires adding the cookies to the request only after all extension passes have finished. The same applies to parsing network requests.
@@
Remark: While BrowserOnly functionality could in principle be incorporated into existing security-relevant attributes (e.g., HttpOnly) merging these functions would lead to changes to the standard cookie usage model. CREAM instead creates a new attribute that can be used independently, preserving existing usage models and backward compatibility.
5.2 Monitored Attribute
@@
The Monitored attribute instructs the browser to maintain a server-defined message and a changelog for the cookie. These values are only accessible to the Network Service. As such, Monitored only requires the Cookie Monster’s Set behavior be updated to log changes to Monitored cookies. As discussed in Section 2.3.2, Set takes a new cookie, checks if a matching cookie already exists, and replaces it if found. Hence when setting a cookie, if a preexisting match is found and it is a Monitored cookie, we check whether the new cookie has a Monitored attribute. If it does not, the matching cookie’s Monitored message and changelog are copied to the new cookie. Then the two cookies are compared and any differences are added to the newly copied changelog. If the new cookie has a Monitored attribute, the new attribute’s message is kept and the changelog is cleared. Similar to BrowserOnly, extensions (and JS) cannot create Monitored cookies. Thus, a cookie with a Monitored attribute must be set by the server and all existing changes must have been seen by the server. This behavior allows the server to update the Monitored attribute as needed and minimize the size of the changelog by removing previously received/validated entries. To avoid resource exhaustion attacks [71], a maximum number of changelog entries can be set. The cookie can simply be flagged as invalid upon reaching the limit.
@@
As discussed in Section 4, along with the Monitored attribute, the browser generates a symmetric key for each site it visits and attaches the key to all outbound HTTPS requests to that site. The browser must maintain a map of keys and their associated sites that is available to URLRequestHTTPJobs. These keys must be created by and maintained within the Network Service. When sending a request, the URLRequestHTTPJob simply fetches the correct key from this map and attaches the key to the request using a custom header. To avoid exposure/tampering by extensions, the key is added after extensions have processed the request headers.
@@
When adding Monitored cookies to network requests (or whenever fetched in general), the changelog and Monitored message are never included in the cookie. As such, Monitored cookies can be seen by extensions. However, when building the Cookie header, we log which cookies are Monitored cookies. Then, after the extensions have processed the headers, we generate the Monitored report for the Monitored cookies. Specifically, for each Monitored cookie, we retrieve the cookie’s changelog, current settings, and Monitored message from the Cookie Monster and add them to the report. Once all the cookies have been added to the report, the report is timestamped and authenticated with the generated key for the site. The report is authenticated as a whole rather than per individual cookie’s sub-report because all cookies of a single site use the same key. Nonetheless, each cookie contains the key within its authenticated/encrypted Monitored message to ensure that the server always receives (and can verify) the key regardless of which set of cookies is present.
@@
Handling Monitored cookies on incoming responses requires no additional changes. Unlike BrowserOnly cookies, Monitored cookies can be seen by extensions. Similarly, even if extensions see the Monitored message, they cannot change it or use it outside of the user’s browser. At best, an extension can remove the attribute and disable the changelog. However, this would only lead the server to reject the cookie as it would not contain the expected report.
6 Security Analysis
@@
We argue that the BrowserOnly and Monitored attributes protect cookies against malicious extensions (and JS), including those with highly privileged API permissions (analyzed in Section 3.1). We argue each attribute’s security separately, as BrowserOnly and Monitored are independent.
6.1 BrowserOnly
@@
$\sf{\mathcal{A}dv}$ can also attempt to access cookies in network headers using the WebRequest and DeclarativeNetRequest APIs. Specifically, $\sf{\mathcal{A}dv}$ can target the Cookie and Set-Cookie HTTP headers. When building network requests, CREAM’s Network Service withholds BrowserOnly cookies from extensions only adding them to the request once all extensions have finished executing. Similarly, when parsing server replies, the Network Service strips any Set-Cookie headers with the BrowserOnly attribute before passing the response to extensions. Therefore, an extension can never access or edit BrowserOnly cookies in network traffic.
@@
6.2 Monitored
Similar to the cases above, $\sf{\mathcal{A}dv}$ may try to get and set cookies using the Cookies API. While Monitored cookies are accessible, the Monitored message and changelog are never returned to the Cookies API. This information is only available through the report generation endpoint internal to the Network Service. Regardless of whether the Monitored message is available, $\sf{\mathcal{A}dv}$ cannot generate a valid report thus invalidating the cookie. $\sf{\mathcal{A}dv}$ does not know the server’s secret key to forge a malicious Monitored message. $\sf{\mathcal{A}dv}$ could try to interleave the Monitored attribute of a different cookie instead, however, the name-value pair in the attribute would not match the stolen cookie. Similarly, $\sf{\mathcal{A}dv}$ cannot change the name-value pair due to the authenticated encryption. Even with the correct Monitored message, $\sf{\mathcal{A}dv}$ would not have the browser’s key and thus cannot generate a valid authenticated Monitored report. $\sf{\mathcal{A}dv}$ also cannot alter the expected key in the Monitored message as they again do not have the server’s secret key.
@@ 6.2 Monitored
$\sf{\mathcal{A}dv}$ could also try to steal and tamper with cookies in network traffic using the WebRequest and DeclarativeNetRequest APIs. Monitored cookies and their messages are accessible through the Cookie and Set-Cookie headers. However, Monitored cookies stolen by $\sf{\mathcal{A}dv}$ are unusable. $\sf{\mathcal{A}dv}$ could also edit Monitored cookies through these headers. Since these edits are done directly to the header, they would not appear in the changelog. As such, editing the Set-Cookie header would alter the cookie’s attributes, but these altered settings would appear in the authenticated Monitored report. As a consequence, the server can see that the cookie’s settings do not match the expected configuration and detect the changes. Similarly, if $\sf{\mathcal{A}dv}$ tried to revert the changes before the report generation, the reverted changes would be logged and detectable. $\sf{\mathcal{A}dv}$ could also remove the Monitored attribute itself. This would disable the changelog, but no report would be generated, making the cookie invalid. $\sf{\mathcal{A}dv}$ can also edit cookies in the Cookie header, but this header only contains the cookie’s name and value. Changing either would invalidate the cookie as it would no longer match the expected name-value pair in the Monitored message. $\sf{\mathcal{A}dv}$ could also attempt to tamper with the Monitored message itself. Attempts to remove, replace, or alter the cookie’s Monitored message would invalidate the cookie as it would either generate no report or an invalid report. Finally, $\sf{\mathcal{A}dv}$ could try to tamper with the key and Monitored report also included in network requests. However both the key and report are attached to the request post extensions’ accesses (recall Section 5) and thus are not accessible to extensions.
@@
$\sf{\mathcal{A}dv}$ could try to access/change the key attached to the network request. However, the browser’s key is only attached to HTTPS traffic. As such, the key is always encrypted by TLS on the network and inaccessible to $\sf{\mathcal{A}dv}$.
7 Implementation & Evaluation
We implement CREAM by extending Chromium version 122.0.6168.0 with 535 lines of C++ code across 31 files. For evaluation, we deploy the prototype on a Ubuntu 20.04 machine running at 3.70 GHz with 32GB of RAM. To test CREAM we implemented two Node.js test servers: an HTTPS web application and a malicious endpoint. The test web application allows a user to register an account and log in. When logging in, the server generates a session cookie with pertinent cookie attributes for each test. To allow the server to create cookies with BrowserOnly and Monitored attributes, we added them to a fork of the Node.js cookie library. The malicious server acts as a relay for any messages sent by our malicious extensions discussed in section 3.1. Our prototype and test network can be found on our open-source repository [39] while our fork of Node.js’s cookie library can be found in a separate repository [72]. We reemphasize that these new attributes are optional and CREAM remains compatible with websites that choose not to use them.
7.1 Cookie Sizes & Browser Storage
Each attribute requires storing additional cookie data. The BrowserOnly attribute adds one byte of memory to each cookie. The Monitored attribute is larger, requiring 49 additional bytes. This is due to the Monitored message and changelog being a string and vector, respectively. Both are 24-byte long classes (in a 64-bit architecture) plus there is an additional byte for the default empty character in memory. Monitored also relies on symmetric keys. CREAM uses Chromium’s built-in AES implementation to create the keys. As such, each key is 256 bits long.
The Monitored message and changelog entries also affect the size of Monitored cookies. Both fields depend on server-defined attributes, namely the cookie’s name, value, domain, and path. Prior studies have found the average length of a cookie’s name-value pair to be 37 characters [73] and its domain to be 14 characters long [74]. We use these averages and a default path of “/” for our analysis. In the worst case, all of a cookie’s attributes change at once and thus are included in the resulting changelog entry. CREAM logs changes in an encoded format (i.e. “SameSite: Strict” as “SS:2”) to reduce the entry’s size. As such, this would result in a 108-byte entry and a total memory cost of 132 bytes including the string object itself. However, in most cases, we expect fewer attributes to change at once leading to lower log entry sizes. The Monitored message’s size also varies on the server’s behavior. Every message contains the cookie’s name, value, and the browser’s key at least. In our tests, the key was 32 bytes long resulting in a 131-byte-long encrypted string.
@@ 7.2 Operation Run-Times
To assess run-time overheads, we measure the time to create, get, and set different cookie types (i.e., “Standard” cookies, Monitored, and BrowserOnly) and compare these times to an unmodified Chromium instance (“Baseline”). This timing depends greatly on the other processes in the browser. As such, we repeat each operation 100,000 times and report the median completion times (Figure 4(C)) to remove the impact of outliers. We measure these times from the Network Service, JS, and extensions. However, creating cookies and all BrowserOnly tests are only measurable from the Network Service. Similarly, all tests are performed synchronously to ensure accuracy.
@@
A cookie’s access time varies on the location of the access. As the Cookie Monster is in the Network Service, accesses from the Network Service are much faster than JS or extensions, which require IPC. The BrowserOnly and Monitored attributes’ additional checks add little overhead to cookie operations. Creating a cookie without either attribute was only 5 $\mu$s slower than the baseline. Fetching a regular cookie from the Network Service, JS, and extensions took 69, 100, and 100 $\mu$s longer than the baseline. Similarly updating a cookie took an additional 100 $\mu$s from both JS and extensions. Curiously, updating a standard cookie through the Network Service was 278 $\mu$s faster than the baseline. Repeating this experiment several times gave us similar results. As CREAM does not optimize cookie operations, this “improvement” reflects the browser’s impact on completion time rather than an actual reduction. Cookie operations in the Network Service are 2$\sim$3 orders of magnitude faster than in JS and extensions (separate processes). As such, operations inside the Network Service are much more impacted by the execution of other processes in the browser likely leading to the discrepancy seen.
@@
BrowserOnly cookies also present a small overhead, requiring 3, 71, and 0 more $\mu$s to be created, retrieved, and updated, respectively. Monitored cookies take an additional 56 $\mu$s to create. Fetching a Monitored cookie through the Network Service, JS, and extensions took an additional 68, 300, and $-100$ $\mu$s. Again the improvement seen when fetching through extensions reflects the impact of other processes and (in this case) IPC on execution time. Updating Monitored cookies took an additional 18, 1,000, and 1,000 $\mu$s respectively. The majority of this overhead stems from updating the cookie’s changelog. Due to this, we also test how varying the numbers of log entries affects a Monitored cookie’s update time. In this test, we timed how long it took to update the cookie 10 times and found the median to be 7,247 $\mu$s. In the original test, updating the cookie once had a median time of 708 $\mu$s. This reflects a linear increase (variability due to background browser operations) indicating that previous log entries do not impact the update time of subsequent entries.
11111.51.51.51.522222.52.52.52.53333BSBOMTime (ms)RequestResponse
(A) Req./Response Handling
000.020.020.020.020.060.060.060.060.10.10.10.1Empty Log1 Entry10 Entries5 CookiesTime (ms)
(B) Report Generation
Figure 5: Latency of (A) different cookie types and (B) Monitored reports for varying number of log entries/cookies.
BO = BrowserOnly; M=Monitored; S=Standard; B=Baseline.
7.3 Server-Side

### v3

@@ V-A BrowserOnly Attribute
Similarly, Figure 3(B) displays how the Network Service parses the server’s response. When the response is received (step B.1 in Figure 3(B)), the URLRequestHTTPJob forwards the response headers to extensions (B.2). Once the extensions finish, the headers are returned to the HTTPJob (B.3). At this point, any new cookies are added to the Cookie Monster (B.4) and the response is sent to the renderer for display (B.5). To hide BrowserOnly cookies in responses, we strip all Set-Cookie headers with the BrowserOnly attribute from the response on arrival (B.1). Then when the headers are returned from the extensions (B.3) we reattach the BrowserOnly headers to the response.
@@
For simplicity, we omit some details of how the Network Service handles transactions and interacts with extensions. Requests pass through several classes between the renderer and the network. We focus on the URLRequestHTTPJob as all BrowserOnly (and Monitored) functionality was added to this class. Similarly, when building a network request, its headers are sent to extensions up to three times [65]. Protecting BrowserOnly cookies requires adding the cookies to the request only after all extension passes have finished. The same applies to parsing network replies.
@@
Remark: While BrowserOnly functionality could in principle be incorporated into existing security-relevant attributes (e.g., HttpOnly) merging these functions would change the standard cookie usage model. CREAM instead creates a new attribute that can be used independently, preserving existing usage models and backward compatibility.
V-B Monitored Attribute
@@
The Monitored attribute instructs the browser to maintain a server-defined message and a changelog for the cookie. These values are only accessible to the Network Service. As such, Monitored only requires updating the Cookie Monster’s Set behavior to log changes made to Monitored cookies. As discussed in Section II-C2, Set takes a new cookie, checks if it already exists, and replaces the existing copy (if found). For Monitored cookies, if a preexisting cookie is found, we check whether the new cookie has a Monitored attribute. If the new cookie does not have the Monitored attribute, the original cookie’s Monitored message and changelog are copied to the new cookie. Then the two cookies are compared and any differences are added to the newly copied changelog. Otherwise, if the new cookie has a Monitored attribute, it replaces the prior Monitored attribute and the changelog is reset. Similar to BrowserOnly, extensions (and JS) cannot create Monitored cookies. Thus only the server can set a cookie with a Monitored attribute. By keeping the new cookie’s attribute the server can update the Monitored attribute as needed. Further, updating the Monitored attribute implies the server has seen the current changelog. Therefore, clearing the changelog removes previously validated entries, minimizing its size. To avoid resource exhaustion attacks [78], a maximum number of changelog entries can be set and a cookie can be flagged as invalid upon reaching this limit.
@@
As discussed in Section IV, along with the Monitored attribute, the browser generates a symmetric key for each site visited and attaches the key to all HTTPS requests to that site. The browser must maintain a map of keys and their associated sites that is available to URLRequestHTTPJobs. These keys must be created by and maintained within the Network Service. When sending a request, the URLRequestHTTPJob simply fetches the correct key from this map and attaches it to the request using a custom header. To avoid exposure/tampering by extensions, the key is added after extensions have processed the request headers.
@@
When adding Monitored cookies to network requests (or whenever fetched in general), the changelog and Monitored message are never included in the cookie. As such, Monitored cookies can be seen by extensions. However, when building the Cookie header, we log which cookies are Monitored cookies. Then, after the extensions have processed the headers, we generate the Monitored report for the Monitored cookies. Specifically, we retrieve each Monitored cookie’s changelog, current settings, and Monitored message from the Cookie Monster and add them to the report. Once all the cookies have been added, the report is timestamped and authenticated with the generated key for the site. The report is authenticated as a whole rather than per cookie’s sub-report as all cookies for a single site use the same key. Nonetheless, each cookie contains the key within its authenticated/encrypted Monitored message to ensure that the server always receives (and can verify) the key regardless of the cookies present.
@@
Handling Monitored cookies on incoming responses requires no additional changes. Unlike BrowserOnly cookies, Monitored cookies can be seen by extensions. Similarly, even if extensions see the Monitored message, they cannot change it or use it outside the user’s browser. At best, an extension can remove the attribute and disable the changelog. However, this would only cause the server to reject the cookie as it would not contain the expected report.
VI Security Analysis
@@
We argue that the BrowserOnly and Monitored attributes protect cookies against malicious extensions, including those with highly privileged API permissions (analyzed in Section III-A). We argue each attribute’s security separately, as BrowserOnly and Monitored are independent.
VI-A BrowserOnly
@@
$\sf{\mathcal{A}dv}$ can also attempt to access cookies in network headers using the WebRequest and DeclarativeNetRequest APIs. Specifically, $\sf{\mathcal{A}dv}$ can target the Cookie and Set-Cookie HTTP headers. When building network requests, CREAM’s Network Service withholds BrowserOnly cookies from the request until after extensions have finished executing. Similarly, when parsing replies, the Network Service strips any Set-Cookie headers with the BrowserOnly attribute before passing the response to extensions. Therefore, an extension can never access or edit BrowserOnly cookies in network traffic.
@@
VI-B Monitored
Similar to the cases above, $\sf{\mathcal{A}dv}$ may try to get and set cookies using the Cookies API. While Monitored cookies are accessible, the Monitored message and changelog are never returned to the Cookies API. This information is only available through the report generation endpoint internal to the Network Service. Regardless of whether the Monitored message is available, $\sf{\mathcal{A}dv}$ cannot generate a valid report thus invalidating the cookie. $\sf{\mathcal{A}dv}$ does not know the server’s secret key to forge a malicious Monitored message. $\sf{\mathcal{A}dv}$ could try to interleave the Monitored attribute of a different cookie instead, however, the name-value pair in the attribute would not match the stolen cookie. Similarly, $\sf{\mathcal{A}dv}$ cannot change the expected name-value pair due to the authenticated encryption. Even with the correct Monitored message, $\sf{\mathcal{A}dv}$ would not have the browser’s key and thus cannot generate a valid authenticated Monitored report. $\sf{\mathcal{A}dv}$ also cannot alter the expected key in the Monitored message as they again do not have the server’s secret key.
@@ VI-B Monitored
$\sf{\mathcal{A}dv}$ could also try to steal and tamper with cookies in network traffic using the WebRequest and DeclarativeNetRequest APIs. Monitored cookies and their messages are accessible through the Cookie and Set-Cookie headers. However, Monitored cookies stolen by $\sf{\mathcal{A}dv}$ are unusable. $\sf{\mathcal{A}dv}$ could also edit Monitored cookies through these headers. Since these edits are done directly to the header, they would not appear in the changelog. As such, editing the Set-Cookie header would alter the cookie’s attributes, but these altered settings would appear in the authenticated Monitored report. As a consequence, the server can see that the cookie’s settings do not match the expected configuration and detect the changes. Similarly, if $\sf{\mathcal{A}dv}$ tried to revert the changes before the report generation, the reverted changes would be logged and detectable. $\sf{\mathcal{A}dv}$ could also remove the Monitored attribute itself. This would disable the changelog, but no report would be generated, making the cookie invalid. $\sf{\mathcal{A}dv}$ can also edit cookies in the Cookie header, but this header only contains the cookie’s name and value. Changing either would invalidate the cookie as it would no longer match the expected name-value pair in the Monitored message. $\sf{\mathcal{A}dv}$ could also attempt to tamper with the Monitored message itself. Attempts to remove, replace, or alter the cookie’s Monitored message would invalidate the cookie as it would either generate no report or an invalid report. Finally, $\sf{\mathcal{A}dv}$ could try to tamper with the key and Monitored report also included in network requests. However both the key and report are attached to the request post extensions’ accesses (recall Section V) and thus are not accessible to extensions.
@@
Finally, $\sf{\mathcal{A}dv}$ could try to access/change the browser’s key attached to network requests. However, these keys are only attached to HTTPS traffic. As such, the key is always encrypted by TLS on the network and inaccessible to $\sf{\mathcal{A}dv}$.
VII Implementation & Evaluation
We implement CREAM by extending Chromium version 122.0.6168.0 with 535 lines of C++ code across 31 files. For evaluation, we deploy our prototype on an Ubuntu 20.04 machine running at 3.70 GHz with 32GB of RAM. To test CREAM we implemented a simple HTTPS web application using Node.js. The test application allows a user to register an account and log in. When logging in, the server generates a session cookie with pertinent cookie attributes for each test. To allow the server to create cookies with BrowserOnly and Monitored attributes, we added them to a fork of the Node.js cookie library. Our prototype and test network can be found on our open-source repository [43] while our fork of Node.js’s cookie library can be found in a separate repository [79]. We reemphasize that usage of these new attributes is optional and CREAM remains fully compatible with websites that choose not to use them.
VII-A Cookie Sizes & Browser Storage
Each attribute requires storing additional cookie data. The BrowserOnly attribute adds one byte of memory to each cookie. The Monitored attribute is larger, requiring 49 additional bytes. This is due to the Monitored message and changelog being a string and vector, respectively. Both are 24-byte long classes (in a 64-bit architecture) plus there is an additional byte for the default empty character in memory. Monitored also relies on symmetric keys. For this, CREAM uses Chromium’s built-in AES implementation which uses 256-bit keys.
The Monitored message and changelog entries also affect the size of Monitored cookies. Both fields depend on server-defined attributes, namely the cookie’s name, value, domain, and path. Prior studies found the average length of a cookie’s name-value pair to be 37 characters [80] and its domain to be 14 characters long [81]. We use these averages and a default path of “/” for our analysis. In the worst case, all of a cookie’s attributes change at once and are included in the resulting changelog entry. CREAM logs changes in an encoded format (i.e. “SameSite: Strict” as “SS:2”) to reduce each entry’s size. As such, this would result in a 108-byte entry and a total memory cost of 132 bytes including the string object itself. However, in most cases, we expect fewer attributes to change at once leading to lower log entry sizes. The Monitored message’s size also varies on the server’s behavior. Every message contains the cookie’s name, value, and the browser’s key at least. In our tests, the key was 32 bytes long resulting in a 131-byte-long Monitored message.
@@ VII-B Operation Run-Times
To assess run-time overheads, we measure the time to create, get, and set different cookie types (i.e., “Standard” cookies, Monitored, and BrowserOnly) and compare them to an unmodified Chromium instance (“Baseline”). This timing depends greatly on the other processes in the browser. As such, we repeat each operation 100,000 times and report the median completion times (Figure 4(C)) to remove the impact of outliers. We measure these times from the Network Service, JS, and extensions as the timing is dependent on where in the browser the operation occurs. As cookies are stored in the Network Service, accesses from the Network Service are generally faster than JS or extensions which require IPC. However, creating cookies and all BrowserOnly tests are only measurable from the Network Service. Similarly, all tests are performed synchronously to ensure accuracy.
@@
CREAM’s additional checks add little overhead to regular cookie operations. Creating cookies without either new attribute was only 5 $\mu$s slower than the baseline. Similarly fetching a cookie from the Network Service, JS, and extensions took 69, 100, and 100 $\mu$s longer than the baseline respectively. Updating a cookie took an additional 100 $\mu$s from both JS and extensions. However, updating a cookie through the Network Service was 278 $\mu$s faster than the baseline. Repeating the experiment several times yielded similar results. As CREAM does not optimize cookie operations, this “improvement” likely reflects the browser’s impact on execution times rather than an actual reduction. The browser is a large multi-process asynchronous system. As such the execution of other processes in the browser (and underlying host) likely led to the discrepancy seen.
@@
BrowserOnly cookies also present little overhead, requiring 3, 71, and 0 more $\mu$s to be created, retrieved, and updated respectively. Monitored cookies took an additional 56 $\mu$s to create. Fetching a Monitored cookie from the Network Service, JS, and extensions took an additional 68, 300, and $-100$ $\mu$s respectively. Again the improvement seen when fetching through extensions reflects the impact of other processes (and IPC in this case) on execution times. Updating Monitored cookies took 18, 1,000, and 1,000 $\mu$s longer than the baseline respectively. Most of this overhead stems from updating the cookie’s changelog. As such, we also test how varying the number of log entries affects a Monitored cookie’s update time. For this, we timed how long it took to update the cookie 10 times and found the median to be 7,247 $\mu$s. In the original test, the median time to update a cookie once was 708 $\mu$s. This change reflects a linear increase (some variability due to background browser operations) indicating previous log entries do not impact the update time of subsequent entries.
VII-C Server-Side
5,00050005{,}0005 , $0006{,}$ $0007{,}$ $0008{,}$ $0001{,}$ $0002{,}$ $0003{,}$ $0004{,}$ $0005{,}$  [^0] Replies/SecondBaselineBrowserOnlyMonitored
Figure 5: Average server throughput (replies/s) for varying concurrent users